OpenStudy (anonymous):

103 is 20% of what number?

OpenStudy (campbell_st):

multiply both by 5 to get 100%

OpenStudy (anonymous):

20% is 103 then 100%=20% *5 = 103 *5

OpenStudy (anonymous):

515 .

OpenStudy (campbell_st):

looks correct to me

OpenStudy (anonymous):

I think its 515

OpenStudy (nincompoop):

look at the pattern 2 is 1/2 of 4 (1/2 is 50%) 2*2 = 4 5 is 1/4 of a number (1/4 is 25%) multiply 5 by 4 103 is 1/5 of a number (1/5 is 20%) 103 * 5

OpenStudy (anonymous):

"103 is 20% of what number?" can be rewritten as the equation below: 103 = 0.20 * x if you wanna make equations from sentences you'll wanna remember that "is" means "=" "%" means "1/100" "what number" means an unknown quantity that we might call "x" or "n" or whatever